A Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 23,000 33,000 56,000
Fixed manufacturing overhead costs $ 730,000 $ 200,000 $ 930,000
Variable manufacturing overhead cost per machine-hour $ 5.50 $ 5.50

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 374,000 $ 330,000 $ 704,000
Direct labor cost $ 210,000 $ 130,000 $ 340,000
Machine-hours 17,000 6,000 23,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 280,000 $ 280,000 $ 560,000
Direct labor cost $ 180,000 $ 280,000 $ 460,000
Machine-hours 6,000 27,000 33,000

The company had no underapplied or overapplied manufacturing overhead during the year.

Required:

1. Assume the compant uses a plantwide predetermined overhead rate based on machine-hours.

a. Compute the plantwide predetermined overhead rate.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If the company establishes bid prices that are 140% of total manufacturing cost, what bid prices would it have established for Job D-70 and Job C-200?

d. What is the companies cost of goods sold for the year?
Answer #1

Answer :-

Assume the compant uses a plantwide predetermined overhead rate based on machine-hours.

1-a. Total Overhead of Molding = Fixed Manufacturing Overhead Cost of Molding + ( Variable manufacturing overhead cost per machine-hours × Machine-hours)

Fixed Manufacturing Overhead Cost of Molding = $730,000

Variable manufacturing overhead cost per machine-hour = $5.50

Machine-hours = 23,000 hours

Total Overhead of Molding = $730,000 + ( $5.50 × 23,000)

Total Overhead of Molding = $856,500

Total Overhead of Fabrication = Fixed Manufacturing Overhead Cost of Fabrication + ( Variable manufacturing overhead cost per machine-hours × Machine-hours)

Fixed Manufacturing Overhead Cost of Fabrication = $200,000

Variable manufacturing overhead cost per machine-hour = $5.50

Machine-hours = 33,000 hours

Total Overhead of Fabrication = $200,000 + ( $5.50 × 33,000)

Total Overhead of Fabrication = $381,500

Total overhead = Total Overhead of Molding + Total Overhead of Fabrication

Total Overhead = $856,500 + $381,500

Total Overhead = $1,238,000

Plantwide Predetermined overhead rate = Total Overhead / Total machine-hours

Total Machine-hours = 56,000 hours

Predetermined Overhead rate = $1,238,000 / 56,000

Predetermined Overhead rate = $22.11 per machine-hour.

1-b.

Job D-70 Job C-200
Direct Materials $704,000 $560,000
Direct Labor $340,000 $460,000
Manufacturing Overhead(Total machine-hours × Predetermined overhead rate) $508,530 (23,000 x 22.11)

$729,630 (33,000 x 22.11)

Total Manufacturing Costs $1,552,530 $1,749,630

1-c.

Bid price of Job D-70 = Total Manufacturing Cost of Job D-70 × 140%

Bid price of Job D-70 = $1,552,530 x 140%

Bid price of Job D-70 = $2,173,542

Bid price of Job C-200 = Total Manufacturing Cost of Job C-200 × 140%

Bid price of Job C-200 = 1,749,630 x 140%

Bid price of Job C-200 = $2,449,482

1-d. Cost of Goods Sold = Total Manufacturing costs of Job D-70 and Job C-200

Cost of good sold = $1,552,530 + $1,749,630

Cost of Good sold = $3,302,160